# Governing Together: Toward Infrastructure for Community-Run Social Media

### v1

@@ 1. Introduction
However, users, content, and interactions typically move across community boundaries. Neither the problems communities deal with nor the governance practices they devise in response are purely endogeneous. For example, a malicious actor creating multiple accounts to spam many communities at once is a broader problem that cannot be addressed by improving the governance of one community alone. Although there is growing empirical evidence of inter-community and ecological dynamics shaping community success (e.g., TeBlunthuis, 2021; Kiene et al., 2025; Hwang et al., 2024; Horta Ribeiro et al., 2025), far less work correspondingly designs systems that would help communities manage and leverage these aspects of how they operate. In this work, we suggest that inter-community governance — mechanisms that intervene in the relationships between communities to shape community governance — is a crucial but undersupported area of design and systems-building that would strengthen community governance and make it more sustainable. Leaders of online communities already find ad hoc ways to learn from one another’s communities (Hwang et al., 2024; Kiene et al., 2025), coordinate on widespread issues (Seering et al., 2022; Matias, 2016), and/or protect themselves from each other (Bono et al., 2024; Jhaver et al., 2018; Wu et al., 2024). At present, social computing systems are rarely designed to support this type of interaction between communities; admins and moderators thus put in extra care and resources into leveraging existing affordances and devising workarounds.
To explore the design space of inter-community governance, we present findings from four design workshops with 24 community organizers and developers in the Fediverse, a decentralized, community-run social media network. On the Fediverse, independently operated and governed communities often have bearing on one another because members of different communities frequently interact, even if those communities have divergent goals, norms, and so on. This results in communities experiencing what Hwang et al. (2025) call governance frictions: incompatibilities in the governance decisions of two communities that undermine the goals of one community in the course of advancing the goals of another. Drawing from workshop discussions about ideal workflows for addressing governance frictions, we present six challenges for coordinating governance in community-run social media around information sharing, managing relationships, and encouraging adoption. The ideas proposed by participants come together to envision an ecosystem of interfaces, infrastructures, and resources that emphasizes three key design principles: modularity, forkability, and polycentricity. We present the challenges, ideas, and principles as conceptual tools that future work can leverage as a road map for designing for inter-community governance in social computing — particularly, for social media.
@@ 2.1. Toward Community-Run Social Media
We define communities as discrete units of collective decision-making within social computing systems (Foote and Hwang, 2023; Bruckman, 2006), where a self-defined group of users freely and voluntarily create and join a shared space toward common interests and have mechanisms to enforce shared values. Community governance thus refers to how these groups self-govern within a system. We focus on community governance in social media systems in particular, which we define per Zhang et al. (2024): “systems that facilitate the sharing of user-generated expressive content, or media, from one user to others.” This includes centralized platform-based systems such as Instagram, Pinterest, and Slack, as well as decentralized protocol-based systems such as Visual IRC, Bluesky, and Mastodon.
@@
Across social media systems, community governance has been a critical part of how governance online is enacted, providing voluntary but significant “civic labor” (Matias, 2019a). Li et al. (2022c) estimates that the monetary value of volunteer moderation by communities on Reddit in 2020 was approximately 3.4 million USD (a conservative estimate). Qualitative work offers rich accounts of the emotional labor (Dosono and Semaan, 2019) and care (Yu et al., 2020) community governance involves, across multiple dimensions of crafting a community space (Morgan et al., 2013; Narayan et al., 2017; Seering et al., 2020). Because social media operates at “unfathomable” scales (Gillespie, 2018), community-level governance can provide nuanced and contextualized decision-making that supports interactions across a much richer array of goals, topics, and values (Chandrasekharan et al., 2018; Weld et al., 2021). In this way, community governance can also be an opportunity for individuals to shape the norms of digital life (Gibson, 2019), nurturing spaces they believe will be valuable (Yu et al., 2020).
@@
Amidst persistent concerns about the outsized power of centralized social media platforms (Douek, 2022; O’Donnell, 2021; DeNardis and Hackl, 2015; Hinds et al., 2020), interest in building, understanding, and strengthening community-owned and -run social media systems has grown (Sahneh et al., 2024; Cohn and Mir, 2022; Mansoux and Abbing, 2020; Masnick, 2019). Community-run social media decentralizes power through shared, open protocols that enable an ecology of interoperable platforms (Masnick, 2019) rather than having a centralized entity that provides oversight and constrains communities. For example, on the Fediverse — a decentralized social media network that is the empirical focus of this work (introduced at length in §3) — communities of users can create or join servers that enable them to interact with others through a shared protocol. Each server self-governs, offering users more control (Cohn and Mir, 2022). However, community governance comes with its own challenges (Hwang et al., 2025; Anaobi et al., 2023). In addition to demanding care and labor (Dosono and Semaan, 2019; Li et al., 2022b), community governance can require knowledge people do not have or even know to anticipate (Tosch et al., 2024). Community governance can thus be difficult to sustain, leading to psychological distress and burnout (Schöpke-Gonzalez et al., 2024).
@@ 2.2.1. Looking Within Communities
Quantitative metrics for improving community governance often measure activity or outcomes within a community, such as growth (Warncke-Wang et al., 2023), turnover (Dabbish et al., 2012), compliance (Matias, 2019b), and toxicity (Jhaver et al., 2021a). Accordingly, researchers and designers have naturally focused on the internal dynamics that appear to lead to desirable outcomes (Kraut and Resnick, 2012). Prior work articulates the practices of moderators, admins and other leaders within a community (Lo, 2018; Seering et al., 2020; Squirrell, 2019; Anaobi et al., 2023), such as their efforts to foster safe spaces (Wu et al., 2024; Dosono and Semaan, 2019; Evans et al., 2018) or onboard newcomers (Ren et al., 2012; Morgan et al., 2013). Related work focuses on understanding the organizational and institutional dynamics of how communities govern in these socio-technical environments (e.g., Jiang et al., 2019; Schneider, 2022; Frey and Sumner, 2019). Researchers have additionally looked to devise and evaluate interventions (e.g., introduction of a new moderation approach) in a community (Tran et al., 2022; TeBlunthuis et al., 2021). Others involve building novel tools that aim to make community governance more sustainable by: reducing the volume of moderation work (Jhaver et al., 2022; Atreja et al., 2024), mediating conflict between members (Doan and Seering, 2025), automatically providing context for decisions (Kuo et al., 2023; Song et al., 2023); and scaffolding policy-making (Zhang et al., 2020).
The literature predominantly seeks to assess dynamics, mechanisms, and interventions within a community (Kraut and Resnick, 2012). Yet, many of the issues these interventions respond to are neither specific nor endogeneous to a particular community. They often reflect ubiquitous organizational problems that communities encounter, such as around technology adoption (Kiene et al., 2019) or improving rule compliance (Matias, 2019b). Other times, they reflect broader problems on the social media, such as a toxicity (Massanari, 2017), harassment campaigns (Han et al., 2023), or other communities acting in bad faith (Kharazian et al., 2023; Heslep and Berge, 2021). In short, communities neither encounter nor address these issues in isolation; examining dynamics within a community alone is valuable but insufficient to address challenges communities face.
@@ 2.2.2. Looking Between Communities
The community ecologies work tends to focus on participation. Related empirical work shows that communities also impact one another’s governance practices. For example, evaluations of deplatforming by platforms from the top-down have found that moderation interventions in one community or system have direct, causal effects on the dynamics of others, albeit with some mixed substantive results (Horta Ribeiro et al., 2023; Mekacher et al., 2023; Ali et al., 2021; Horta Ribeiro et al., 2023). Even when communities don’t directly engage with one another, they can experience spillover effects (Russo et al., 2024). Subsequently, communities may want to sanction one another (e.g., community-level blocking), which can reduce activity in the sanctioned community (Colglazier et al., 2024). Conversely, communities may want to cooperate or collaborate on issues, such as sharing strategies, setting boundaries, or coordinating against a bad actor. Research already documents how communities look to one another for governance issues. For example, recent work by Kiene et al. (2025) describes how online communities devise their rules in fundamentally relational ways: looking to other communities that they copy from, distinguish themselves from, and assert their boundaries against.
@@ 2.3.1. Institutional Design
We draw inspiration from work investigating how communities engage in institutional design, devising incentive systems that guide human behavior (North, 1991). Prior work both draws on concepts from institutional theory to recommend new layers of governing to design for (Frey et al., 2019) and builds tools that would enable users to engage in decision-making about governance (Wang et al., 2024; Zhang et al., 2020). For example, Zhang et al. (2020) presents PolicyKit, a software infrastructure that allows members of a community to author and implement new governance procedures on their home systems over time. Although these studies do not focus on inter-community governance, they suggest the potential for developing extensions, tools, and interfaces that might help communities thoughtfully shape interactions they have with other communities. Considering institutional design as an inspiration for designing inter-community governance mechanisms also suggests expanding the scope of design beyond the technical: social rules, policies and incentives, and technical tools all become intertwined levers that can be “designed” to affect and implement inter-community governance.
@@ 2.3.2. Building Community Capacities
Work across HCI focused on how collective action informs our view of how to support coordination in community-run social media. Notions of “social design” from Le Dantec (2016) — which emphasizes designing for social action with communities — and “infrastructuring” community from Ehn (2008) both emphasize building community capacities to act and organize (e.g., Li et al., 2018; Salehi et al., 2015; Irani and Silberman, 2013). We similarly focus on how community capacities for inter-community governance can be expanded.
@@ 3. Empirical Setting
The Fediverse presents an ideal setting to study inter-community governance for two main reasons. First, communities are the main level of decision-making in this context. Although the software a community uses (e.g., Mastodon) can constrain it, communities are largely autonomous in that no central entity prescribes norms or practices to them and they can choose the suite of technologies they rely on. Thus, no centralized entity might step in to address frictions for communities. Second, because users across communities frequently interact, governance frictions regularly become apparent around incidents such as norm and privacy violations (Hwang et al., 2025). Although the protocol provides one sanctioning mechanism – defederation — it provides no mechanisms for communities to proactively identify frictions, resolve disputes, or coordinate actions. Both of these make the importance of designing for inter-community governance particularly critical in this context.
@@ 4.1. Design Workshops and Follow-up Interviews
Figure 1. Guided flow of discussion during each workshop across two main break-out group sessions, followed by optional follow-up interview calls with individuals who chose to take part.:
@@
Table 1. Participants across the workshop. Highlighted indicates follow-up interview.:
We conducted four design workshops running two to three hours over Zoom throughout June and July 2025 with small groups of individuals in the Fediverse, with activities scaffolded with a collaborative virtual whiteboard on Miro. Participants were offered an optional \$60 via Zelle as thanks for their time. The study was reviewed and approved by Princeton University and Northwestern University.
@@
Following introductions, each workshop began with a discussion of an example of cooperation among competing groups of lobster fishermen in Maine, inspired by the ethnographic work of Acheson (1988). This served as a prompt and warm-up activity for participants (Harrington and Dillahunt, 2021; Andersen and Wakkary, 2019), to help them start thinking about themes relevant to governance. Participants were asked to share wonderings and reflections about the case, particularly in comparison to the Fediverse. The workshop then moved into its main activities, the flow of which is captured in Figure 1. First, participants were asked to collaboratively map out key community decisions, before selecting 1-3 that they found the most important. Then, participants took five minutes to sketch out a scenario where a governance friction (Hwang et al., 2025) around any of the decisions arises. We leaned on this concept because it evokes the idea of design frictions, which typically refers to difficulties in user experience. Insofar that differences in how communities govern can cause challenges for one or the other, such differences cause friction. However, as Cox et al. (2016) notes, frictions can also give pause so as to create more mindful interaction. Likewise, we approach governance frictions as a generative opportunity.
@@
Finally, participants were guided through a discussion to brainstorm ideal ways to address the friction they had identified, ideating potential new tools that might support them. Participants were told tools need not be technical in nature: they could be resources, organizations, processes, and so on. Here, we drew inspiration from speculative design approaches (Dunne and Raby, 2013; Coulton et al., 2017; Wong et al., 2017), which emphasize that people can imagine possible worlds by not only coming up with new technologies, but also by imagining alternative social values, organizations, rules, policies, and infrastructures. We found this perspective to be useful when considering the design of governance systems, which often require thinking about design at an institutional level, beyond technical tools alone.
@@
We also conducted 13 optional follow-up interviews in July 2025 (this did not lead to more compensation); participants who completed these are indicated by the rows in blue in Table 1. Interviews ran between 25-56 minutes each and were recorded, with the exception of one (P6). These calls enabled us to ask participants their reflections on the ideas discussed in the workshop sessions. Some participants were also involved in ongoing projects relevant to the issues discussed in workshops; conversations with these individuals also focused on the motivations and goals of their projects.

### v2

@@ 1. Introduction
However, users, content, and interactions typically move across community boundaries. Neither the problems communities deal with nor the governance practices they devise in response are purely endogeneous. For example, a malicious actor creating multiple accounts to spam many communities at once is a broader problem that cannot be addressed by improving the governance of one community alone. Although there is growing empirical evidence of inter-community and ecological dynamics shaping community success (e.g., TeBlunthuis, 2021; Kiene et al., 2025; Horta Ribeiro et al., 2025), far less work designs systems that would help communities manage and leverage these aspects of how they operate. In this work, we suggest that inter-community governance — mechanisms that intervene in the relationships between communities to shape community governance — is a crucial but undersupported area of design and systems-building that would strengthen community governance and make it more sustainable. Leaders of online communities already find ad hoc ways to learn from one another’s communities (Kiene et al., 2025), coordinate on widespread issues (Seering et al., 2022; Matias, 2016), and/or protect themselves from each other (Bono et al., 2024; Jhaver et al., 2018; Wu et al., 2024). At present, social computing systems are rarely designed to support this type of interaction between communities; admins and moderators thus put extra care and resources into leveraging existing affordances and devising workarounds.
To explore the design space of inter-community governance, we present findings from four design workshops with 24 community organizers and developers in the Fediverse, a decentralized, community-run social media network. On the Fediverse, independently operated and governed communities often have bearing on one another because members of different communities interact, even if those communities have divergent goals, norms, and so on. This results in communities frequently experiencing what Hwang et al. (2025) call governance frictions: incompatibilities in the governance decisions of two communities that undermine the goals of one community in the course of advancing the goals of another. Drawing from workshop discussions about ideal workflows for addressing governance frictions, we identify six challenges for coordinating governance in community-run social media around information sharing, managing relationships, and encouraging adoption. The ideas proposed by participants envision an ecosystem of interfaces, infrastructures, and resources that emphasizes three key design principles: modularity, forkability, and polycentricity. We present the challenges, ideas, and principles as conceptual tools that future work can leverage as a road map for designing for inter-community governance in social computing.
@@ 2.1. Toward Community-Run Social Media
We define communities as discrete units of collective decision-making within social computing systems (Foote and Hwang, 2023; Bruckman, 2006), where a self-defined group of users freely and voluntarily create and join a shared space toward common interests and have mechanisms to enforce shared values. Community governance thus refers to how these groups self-govern within a system. We focus on community governance in social media systems in particular, which we define per Zhang et al. (2024): “systems that facilitate the sharing of user-generated expressive content, or media, from one user to others.” This includes centralized platform systems such as Instagram, Pinterest, and Slack, as well as decentralized protocol-based systems such as Visual IRC, Bluesky, and Mastodon.
@@
Across social media systems, community governance has been a critical part of how governance online is enacted, providing voluntary but significant “civic labor” (Matias, 2019b). Li et al. (2022c) estimate that the monetary value of volunteer moderation by communities on Reddit in 2020 was approximately 3.4 million USD (a conservative estimate). Qualitative work offers rich accounts of the emotional labor (Dosono and Semaan, 2019) and care (Yu et al., 2020) community governance involves, across multiple dimensions of crafting a community space (Morgan et al., 2013; Narayan et al., 2017; Seering et al., 2020). Because social media operates at “unfathomable” scales (Gillespie, 2018), community-level governance can provide nuanced and contextualized decision-making that supports interactions across a much richer array of goals, topics, and values (Chandrasekharan et al., 2018; Weld et al., 2021). In this way, community governance can also be an opportunity for individuals to shape the norms of digital life (Gibson, 2019), nurturing spaces they find valuable (Yu et al., 2020).
@@
Amidst persistent concerns about the outsized power of centralized social media platforms (Douek, 2022; O’Donnell, 2021; DeNardis and Hackl, 2015; Hinds et al., 2020), interest in building, understanding, and strengthening community-owned and -run social media systems has grown (Sahneh et al., 2024; Cohn and Mir, 2022; Mansoux and Abbing, 2020; Masnick, 2019). Community-run social media decentralizes power through shared, open protocols that enable an ecology of interoperable platforms (Masnick, 2019) rather than a centralized entity that provides oversight and constrains communities. For example, on the Fediverse — a decentralized social media network that is the empirical focus of this work (introduced at length in §3) — communities of users can create or join servers that enable them to interact with others through a shared protocol. Each server self-governs, offering users more control (Cohn and Mir, 2022). However, community governance comes with its own challenges (Hwang et al., 2025; Anaobi et al., 2023). In addition to demanding care and labor (Dosono and Semaan, 2019; Li et al., 2022b), community governance can require knowledge people do not have or even know to anticipate (Tosch et al., 2024). Community governance can thus be difficult to sustain, leading to psychological distress and burnout (Schöpke-Gonzalez et al., 2024).
@@ 2.2.1. Looking Within Communities
Quantitative metrics for improving community governance often measure activity or outcomes within a community, such as growth (Warncke-Wang et al., 2023), turnover (Dabbish et al., 2012), compliance (Matias, 2019a), and toxicity (Jhaver et al., 2021a). Accordingly, researchers and designers have focused on the internal dynamics that appear to lead to desirable outcomes (Kraut and Resnick, 2012). Prior work articulates the practices of moderators, admins and other leaders within a community (Lo, 2018; Seering et al., 2020; Squirrell, 2019; Anaobi et al., 2023), such as their efforts to foster safe spaces (Wu et al., 2024; Dosono and Semaan, 2019; Evans et al., 2018) or onboard newcomers (Ren et al., 2012; Morgan et al., 2013). Related work focuses on understanding the organizational and institutional dynamics of how communities govern in these socio-technical environments (e.g., Jiang et al., 2019; Schneider, 2022; Frey and Sumner, 2019). Researchers have additionally looked to devise and evaluate interventions (e.g., introduction of a new moderation approach) in a community (Tran et al., 2022; TeBlunthuis et al., 2021). Others involve building novel tools that aim to make community governance more sustainable by: reducing the volume of moderation work (Jhaver et al., 2022; Atreja et al., 2024), mediating conflict between members (Doan and Seering, 2025), automatically providing context for decisions (Kuo et al., 2023; Song et al., 2023); and scaffolding policy-making (Zhang et al., 2020).
The literature predominantly seeks to assess dynamics, mechanisms, and interventions within a community (Kraut and Resnick, 2012). Yet, many of the issues these interventions respond to are neither specific nor endogeneous to a particular community. They often reflect ubiquitous organizational problems that communities encounter, such as around technology adoption (Kiene et al., 2019) or improving rule compliance (Matias, 2019a). Other times, they reflect broader problems on the social media, such as toxicity (Massanari, 2017), harassment campaigns (Han et al., 2023), or other communities acting in bad faith (Kharazian et al., 2023; Heslep and family=Berge, 2021). In short, communities neither encounter nor address these issues in isolation; examining dynamics within a community alone is valuable but insufficient to address challenges communities face.
@@ 2.2.2. Looking Between Communities
The community ecologies work tends to focus on participation. Related empirical work shows that communities also impact one another’s governance practices. For example, evaluations of deplatforming by platforms from the top-down have found that moderation interventions in one community or system have direct, causal effects on the dynamics of others, albeit with some mixed substantive results (Horta Ribeiro et al., 2023; Mekacher et al., 2023; Ali et al., 2021; Horta Ribeiro et al., 2023). Even when communities don’t directly engage with one another, they can experience spillover effects (Russo et al., 2024). Communities may also sanction one another (e.g., community-level blocking), which can reduce activity in the sanctioned community (Colglazier et al., 2024). Conversely, communities may want to cooperate or collaborate on issues, such as sharing strategies, setting boundaries, or coordinating against a bad actor. Research already documents how communities look to one another for governance issues. For example, recent work by Kiene et al. (2025) describes how online communities devise their rules in relational ways: looking to other communities that they copy from, distinguish themselves from, and assert their boundaries against.
@@ 2.3.1. Institutional Design
We draw inspiration from work investigating how communities engage in institutional design, devising incentive systems that guide human behavior (North, 1991). Prior work both draws on concepts from institutional theory to recommend new layers of governing to design for (Frey et al., 2019; Schneider et al., 2021) and builds tools that would enable users to engage in decision-making about governance (Wang et al., 2024; Zhang et al., 2020). For example, Zhang et al. (2020) presents PolicyKit, a software infrastructure that allows members of a community to author and implement new governance procedures on their home systems over time. Although these studies do not focus on inter-community governance, they suggest developing extensions, tools, and interfaces might help communities thoughtfully shape interactions they have with other communities. Considering institutional design as an inspiration for designing inter-community governance mechanisms also suggests expanding the scope of design beyond the technical: social rules, policies and incentives, and technical tools all become intertwined levers that can be “designed” to affect and implement inter-community governance.
@@ 2.3.2. Building Community Capacities
Work across HCI focused on how collective action informs our view of how to support coordination. Notions of “social design” from Le Dantec (2016) — which emphasizes designing for social action with communities — and “infrastructuring” community from Ehn (2008) both emphasize building community capacities to act and organize (e.g., Li et al., 2018; Salehi et al., 2015; Irani and Silberman, 2013). Recent work has proposed affordances and tools that might enable collective and communal forms of governance more generally (Mahar et al., 2018; Jahanbakhsh et al., 2022), such as leaning on peers to help filter email harassment in Squadbox (Mahar et al., 2018).
@@ 3. Empirical Setting
While inter-community governance is relevant for any context with multiple communities, the Fediverse presents an ideal setting to explore the concept for two main reasons. First, communities are the main level of decision-making in this context. Although the software a community uses (e.g., Mastodon) can constrain it, communities are largely autonomous in that no central entity prescribes norms or practices to them and they can choose the suite of technologies they rely on. Thus, no centralized entity might step in to address frictions for communities. Second, because users across communities frequently interact, governance frictions regularly become apparent around incidents such as norm and privacy violations (Hwang et al., 2025). Although the protocol provides one sanctioning mechanism — defederation — it provides no mechanisms for communities to proactively identify frictions, resolve disputes, or coordinate actions. Both of these make the importance of designing for inter-community governance particularly critical in this context.
@@ 4.1. Design Workshops and Follow-up Interviews
Table 1. Participants across the workshop. Highlighted indicates follow-up interview.
The table provides description of the participants, by columns: Workshop number, Participant ID, Age (in brackets), gender, (M, F, or NB), Country, Main Fediverse Role(s), and Occupation. In Workshop #1: P1 (Age 55–64, Non-binary, USA): Roles – Admin, Writer; Occupation – Software Engineer. P2 (Age 18–24, Male, USA): Roles – Admin; Occupation – Graduate Student. P3 (Age 25–34, Male, USA): Roles – User, Researcher; Occupation – Operations Manager. P4 (Age 55–64, Male, USA): Roles – Moderator, Organizer; Occupation – Co-op Organizer & Educator. In Workshop #2: P5 (Age 45–54, Male, USA): Roles – Admin, Organizer; Occupation – Non-profit Management. P6 (Age 35–44, Male, Italy): Roles – Developer; Occupation – Coder. P7 (Age 25–34, Female, UK): Roles – Organizer; Occupation – Head of Social & Community. P8 (Age 35–44, Non-binary, Spain): Roles – Moderator, Organizer; Occupation – Freelance Artist & Educator. P9 (Age 45–54, Male, USA): Roles – Admin; Occupation – Writer. P10 (Age 35–44, Male, Canada): Roles – Moderator, Developer; Occupation – Web Developer. In Workshop #3: P11 (Age 35–44, Male, USA): Roles – Admin, Researcher, Developer; Occupation – Computer Programmer. P12 (Age 25–34, Male, UK): Roles – Organizer; Occupation – Foundation Ambassador. P13 (Age 35–44, Female, Netherlands): Roles – Admin, Organizer; Occupation – OSS Trust & Safety. P14 (Age 35–44, Male, Netherlands): Roles – Writer; Occupation – Writer. P15 (Age 35–44, Non-binary, UK): Roles – Developer; Occupation – Software Engineer. In Workshop #4: P16 (Age 25–34, Male, USA): Roles – Admin; Occupation – Cyber Security. P17 (Age 35–44, Female, USA): Roles – Developer; Occupation – Software Engineer. P18 (Age 18–24, Male, USA): Roles – User, Researcher; Occupation – Graduate Student. P19 (Age 25–34, Male, Czechia): Roles – Admin, Developer; Occupation – Software Architect. P20 (Age 35–44, Non-binary, USA): Roles – Admin; Occupation – Software Engineer. P21 (Age 55–64, Male, Argentina): Roles – Moderator, Organizer; Occupation – Researcher & Designer. P22 (Age 25–34, Male, Brazil): Roles – Moderator; Occupation – Web Developer. P23 (Age 45–54, Female, USA): Roles – Moderator; Occupation – Student. P24 (Age 45–54, Male, USA): Roles – User, Developer; Occupation – Software Engineer.
Figure 1. A screenshot of the slide shown to participants to present the warm-up example of lobster fisherman, which the workshop facilitator talked through.
The figure shows a screenshot of a slide from the workshop, which describes an example that served as the warm-up activity. The text reads: In the 80s, lobster fishermen in Maine identified with “harbor gangs,” informal groups of fishers associated with specific coves or harbors. Fishers want to defend their gang’s territory, as well as make sure that everyone is fishing fairly and sustainably so the seasons to come may be plentiful. In bullet points: Clear markers of which cooperative a fisher is from (color-coded gear); Informal understanding that violators can be sanctioned: cutting fishing lines and traps that don’t respect territorial lines; Widespread, collective norm to sanction of freeloaders; Lobbying regulators together + sharing information e.g., weather or warnings.
We conducted four design workshops (June-July 2025), each running two to three hours over Zoom with small groups of individuals in the Fediverse. Discussions were scaffolded with a collaborative virtual whiteboard on Miro. Participants were offered an optional \$60 via Zelle as thanks for their time. The study was reviewed and approved by [anonymized university ethics review board(s)].
@@
Following introductions, each workshop began with an example of cooperation among competing groups of lobster fishermen in Maine, inspired by the ethnographic work of Acheson (1988). This example — which describes how fishermen manage commons-based resources together for long-term, economic viability of their livelihoods as a whole (Figure 1) — served as a prompt and warm-up activity for participants (Harrington and Dillahunt, 2021; Andersen and Wakkary, 2019). The focus on sustainability and collective benefit in the example oriented the tone of what values guided the workshop session. Participants were asked to share wonderings and reflections about the case, particularly in comparison to the Fediverse. For example, participants noted how the local, in-person nature of the case often created social norms and expectations that incentivized cooperation and made sanctions effective.
Figure 2. Guided flow of discussion during each workshop across two main break-out group sessions, followed by opt-in follow-up interview calls with individuals.
The figure shows a flow of the main workshop activities: (1) Map out key community decisions, which participants then narrow in on one focal decision; (2) Sketch out an issue or scenario where a governance friction around a decision arises; (3) Discuss ideal workflows to address the friction, imagining new tools, organizations, and processes to support these workflows. Finally, there are optional follow-up calls to reflection on workshop discussion.
@@
The workshop then moved into its main activities, the flow of which is captured in Figure 2 (with details in Appendix A). Participants were first asked to collaboratively map out key community decisions, before selecting 1-3 that they found the most important. This was open-ended and we did not provide specific examples, encouraging participants to spitball. Then, participants took five minutes to sketch out a scenario where a governance friction (Hwang et al., 2025) around any of the decisions arises. We leaned on this concept because it evokes the idea of design frictions, which typically refers to difficulties in user experience. Insofar that differences in how communities govern can cause challenges for one or the other, such differences cause friction. However, as Cox et al. (2016) notes, frictions can also prompt more mindful interaction. Likewise, we approach governance frictions as a generative opportunity.
@@
Participants were then asked to brainstorm ideal ways to address the friction they had identified, ideating potential new tools that support them. Participants were told tools need not be technical in nature: they could be resources, organizations, processes, and so on. Here, we drew inspiration from speculative design approaches (Dunne and Raby, 2013; Coulton et al., 2017; Wong et al., 2017), which emphasize that people can imagine possible worlds by not only coming up with new technologies, but also by imagining alternative social values, organizations, rules, policies, and infrastructures. We found this perspective to be useful when considering the design of governance systems, which often require thinking about design at an institutional level, beyond technical tools alone.
@@
Across the activities, we served primarily as note-takers on the Miro, checking in with participants that notes captured were accurate. If conversation lulled, we also prompted participants for their thoughts to restart discussion. Finally, we conducted 13 optional follow-up interviews in July 2025 (this did not lead to more compensation); participants who completed these are indicated by the rows in blue in Table 1. Interviews ran between 25-56 minutes each and were recorded, with the exception of one (P6). These calls enabled us to ask participants their reflections on the ideas discussed in the workshop sessions. Some participants were also involved in ongoing projects relevant to the issues discussed in workshops; conversations with these individuals also focused on the motivations and goals of their projects.